# Spin-free orbital entropy, mutual information, and entanglement analysis

### v1

@@ 4.1 A non-interacting dimer of CH2 diradicals
Table 1 lists the orbital entropies computed from the FCI wave functions of the CH2 diradical and its dimers, for different values of the spin square operator $S^{2}$ and spin projection $M_{s}$. The spin-free entropies are given first, with the “traditional” spin-including ones following in parentheses. The most interesting are orbitals 3 and 4, which form the open shells of the monomer’s triplet state. It can be seen, that already for the monomer, the spin-free orbital entropies are invariant to the choice of the $M_{s}$ component of the triplet state, while the “traditional” entropies change their values. In particular, the spin coupling of HOMO and LUMO orbitals dramatically increases their entropy with respect to the spin-free one. Since we employed a determinant-based, not spin-adapted FCI code, for the dimer it did not always converge to an eigenfunction of $S^{2}$, but to an arbitrary linear combination of the degenerate eigenfunctions. By using a random initial guess, we achieved a fractional value of $\langle S^{2}\rangle$, yet the spin-free entropies were not affected. Moreover, for a spin-unpure state of the dimer the “traditional” entropies can differ for the two degenerate orbitals belonging to individual monomers, while the spin-free ones are the same.
@@
System $\langle S^{2}\rangle$, $M_{s}$ Orbital entropies
@@
For the mutual information, there are too many data to show them all in a table, but it again turns out that the spin-free mutual information is invariant to the choice of spin state of the noninteracting dimer and is (up to a permutation of degenerate orbitals) a direct sum of the mutual information matrices of the monomers, while the “traditional” mutual information does not have these properties. The most interesting is the 4 x 4 submatrix corresponding to the two pairs of open shell orbitals of the monomers. The spin-free mutual information in this subspace vanishes, while the “traditional” one can acquire large values up to 0.31 for some spin couplings. It is thus evident that the spin-free entanglement analysis has some advantages, being useful at least as a supplementary to the “traditional” one.
@@ 4.2 A complex with one iron atom
As a first realistic example of the presented spin-free entanglement analysis we turn to the [Fe(SCH3)4]- complex, which contains an Fe(III) center in a S4 symmetry coordination environment. The ground state follows the Hund’s rule and turns to be a sextet, with each of the five MOs formed from 3d-orbitals of iron being singly occupied. The DMRG calculations of the sextet for the three values of $M_{s}$ = 1/2, 3/2, and 5/2 yield energy -3011.13544(6-8) a.u., with differences in the order of micro Hartrees. Reconstruction of the leading terms of the wave function from the MPS form confirmed that the $M_{s}=5/2$ component is single-determinantal, while the $M_{s}=3/2$ component is dominated by 5 equally contributing Slater determinants, where four of the MOs corresponding to the iron 3d-orbitals have $\alpha$-spin and one has $\beta$ spin. For $M_{s}=1/2$, (${5\choose 2}=10$ Slater determinants with three $\alpha$ and two $\beta$ spins contribute equally.
Figure 2: Entanglement analysis for the ground state ($S=5/2$) of the [Fe(SCH3)4]- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well. The top row contains results for $M_{s}=5/2$ , middle row corresponds to $M_{s}=3/2$ and the bottom one has $M_{s}=1/2$. In the entanglement graph, the orbitals in their ascending energy order are placed in the clockwise direction starting from the gray indicated “noon”.
@@
Figure 2 shows the entanglement analysis for this state and its three positive $M_{s}$ components, in the spin-dependent and spin-free forms. It can be seen that the three spin-free subfigures in the right column are identical, demonstrating the invariance of the spin-free entropy and mutual information with respect to the $M_{s}$ value. The top left subfigure corresponds to the spin-dependent analysis for the $M_{s}=5/2$ component, which is dominated by a single Slater determinant, and thus yield virtually identical results to the spin-free analysis. On the other hand, the middle and bottom left subfigures for $M_{s}=3/2$ and $M_{s}=1/2$, respectively, show strong orbital entropies for the MOs formed from 3d-orbitals of iron, and strong mutual information between each pair of them. For $M_{s}=1/2$ the values are somewhat bigger, in line with 10 Slater determinants contributing rather than five.
@@
The approach can be used for excited states as well. For example, the first excited state with energy -3011.0279(49-50) a.u. is a quartet, with its $M_{s}=3/2$ component dominated by the determinant $2\alpha\alpha\alpha 0$ with a non-negligible contribution of $0\alpha\alpha\alpha 2$, while the $M_{s}=1/2$ component has equal dominant contribution of determinants $2\beta\alpha\alpha 0$, $2\alpha\beta\alpha 0$, and $2\alpha\alpha\beta 0$ with a non-negligible contribution of analogous $0\ldots 2$ determinants. Figure 3 shows the corresponding entanglement analysis for the two positive $M_{s}$ components of this state. It can again be seen that the spin-free entropies and mutual informations are invariant to $M_{s}$, and they agree with the high-spin spin-dependent case, showing strong entropies and mutual information for the first and last of the iron-3d dominated MOs. The entanglement analysis for the $M_{s}=1/2$ component in the spin-dependent form shows in addition to this strong entropies and mutual information between the remaining three iron-3d dominated MOs, which are purely a consequence of the spin coupling.
@@
Figure 4 shows the corresponding entanglement analysis for the second excited state, which is a doublet with energy -3011.016680 a.u.. Its wave function is dominated by several triples of equally contributing Slater determinants, where one of the 5 iron 3d MOs is doubly occupied, one empty, and the remaining three have two $\alpha$ and one $\beta$ electrons, in all three possible orderings. The static correlation due to spin couplings is thus inherently mixed with the strong non-dynamic correlation. This is reflected in the graphical analysis by the simplification of the mutual information graph when going from the spin-dependent to the spin-free analysis, and somewhat lower spin-free orbital entropies, particularly for one of the five orbitals.
@@
Table 2: Spin-free and spin-including total entropies $\tilde{S}_{\rm tot}$, $S_{\rm tot}$ and mutual informations $\tilde{I}_{\rm tot}$, $I_{\rm tot}$ for three states of the [Fe(SCH3)4]- complex and their $M_{s}$ components.
@@
5/2 5/2 0.085 0.087 0.022 0.041
5/2 3/2 0.085 2.588 0.022 1.677
5/2 1/2 0.085 3.451 0.022 2.278
3/2 3/2 0.879 0.891 0.449 0.541
3/2 1/2 0.879 2.792 0.449 1.480
1/2 1/2 3.709 5.623 1.104 2.186
An overview of the spin-free and spin-including total entropies and mutual informations for the three states of this complex are given in Tab. 2. As can be seen, the spin-free quantities are invariant with respect to $M_{s}$, while the spin-including ones strongly depend on the spin projection, with a growing trend towards smaller $M_{s}$. This can be explained by the increased number of Slater determinants forming low-$M_{s}$ components of the spin eigenfunctions, resulting in increased static correlation. However, even the spin-free total entropy and mutual information have a growing trend with decreasing spin $S$, indicating more correlation present in the lower-spin states. This trend is somewhat hidden in the spin-including correlation measures due to their fluctuation with $M_{s}$, so in our view the spin-free quantities are a useful tool to uncover this dependence.
We have thus seen that for states of all spin multiplicities, the spin-free entanglement analysis was able to simplify the picture by filtering out the static correlation due to spin couplings. In particular, the spin-dependent mutual information graph looks almost identical for the $M_{s}=1/2$ component of the doublet and sextet states, while their spin-free graphs are dramatically different, reflecting the different origin of the correlation. It thus seems to be useful to accompany the “traditional” entanglement analysis with its spin-free version.
Figure 3: Entanglement analysis for the first excited state ($S=3/2$) of the [Fe(SCH3)4]- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well. The top row contains results for $M_{s}=3/2$, while the bottom one has $M_{s}=1/2$.
@@
Figure 4: Entanglement analysis for the second excited state ($S=1/2$) of the [Fe(SCH3)4]- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well.
@@ 4.3 A complex with two iron atoms
The [Fe2S2(SCH3)4]2- complex is a model of the biologically important iron-sulfur clusters, which act as catalytic centers [18]. This species exhibits strong electronic correlation due to the coupling of the 3d-electrons of the two iron atoms, while the dynamic correlation is very important to get the correct energetic order of the states of different spin multiplicities (i.e. singlet ground state) [16]. In this work we did not aim to perform high-accuracy calculations of the energies of the states, rather we used the complex as a non-trivial example for the entanglement analysis. Table 3 shows the energies of the $S=0-5$ states if this complex, computed in two different DMRG active spaces. As can be seen, DMRG with a very small active space like e.g. (14,14) yields inverted energy order of the states, however, DMRG(22,21) already yields a qualitatively correct ordering, although the excitation energies are substantially underestimated with respect to DMRG-AC0 [16] due to the lack of dynamic correlation.
@@
Figure 5: Entanglement analysis for the ground ($S=0$) state of the [Fe2S2(SCH3)4]2- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well.
@@
Fig. 5 shows the entanglement analysis for the ground ($S=0$) state. It clearly shows the presence of the full pairwise “anti-ferromagnetic” coupling between the pairs formed from the 3d-dominated MOs of the two iron atoms. Smaller contributions in the wave function come from determinants with some open shells, which have to spin-couple to singlets, and this is reflected by the weak mutual information lines, which disappear in the spin-free subfigure on the right, and the slightly lower spin-free orbital entropies.
Figure 6: Entanglement analysis for the first excited ($S=1$) state of the [Fe2S2(SCH3)4]2- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well. The $M_{s}=1$ component is in the top row, with $M_{s}=0$ component below.
@@
Fig. 6 shows the results for the first excited ($S=1$) state in its two $M_{s}=1,0$ components. The overall situation is similar to the ground state, except that here one pair of orbitals remains open shell “ferromagnetically” triplet-coupled. The spin-including mutual informations differ for the two $M_{s}$ components, while the spin-free ones are identical and a similar “weakening” of the mutual information and orbital entropies as in the ground state happens. The situation is analogous for the other excited states, with the most dramatic spin-free “simplification” occurring for the $M_{s}<S$ components of the “totally ferromagnetic” $S=5$ excited state (analogous to the $S=5/2$ state of the single-iron complex, cf. Fig. 2). Rather than showing figures for all the cases, we summarize the results in Tab. 4. The somewhat less than perfect $M_{s}$-invariance of the spin-free quantities for higher DMRG roots corresponds to the truncation of the bond dimension, which was limited to 2048 for all states and is more limiting to the low-$M_{s}$ states being higher roots. In some cases we did not achieve convergence at all for the higher roots. Nevertheless, the total spin-free entropy $\tilde{S}_{\rm tot}$ shows the same increasing trend towards low spin states as in the complex with single iron atom. On the other hand, the total spin-free mutual information $\tilde{I}_{\rm tot}$ is very similar (around 2.2) for states with $S=0-4$ and decreases to 0.13 only for the highest spin state $S=5$. This is in line with the observation from Figs. 5,6, which show strong pairwise entanglement between orbitals from the two iron atoms, both for the ground state $(S=0)$ and first excited state $(S=1)$. This entanglement “survives” in the spin-free analysis and is thus not caused by spin coupling, but reflects the inherent strong correlation in the system.
Table 4: Spin-free and spin-including total entropies $\tilde{S}_{\rm tot}$, $S_{\rm tot}$ and mutual informations $\tilde{I}_{\rm tot}$, $I_{\rm tot}$ for three states of the [Fe2S2(SCH3)4]2- complex and their $M_{s}$ components.

### v2

@@ 4.1 A non-interacting dimer of CH2 diradicals
Table 1 lists the orbital entropies computed from the FCI wave functions of the CH2 diradical and its dimers, for different values of the spin square operator $S^{2}$ and spin projection $M_{s}$. The spin-free entropies are given first, with the “traditional” spin-including ones following in parentheses. Figure 2 shows the graphical representation of the orbitals. The most interesting are the open shell (active) orbitals (3 and 4), which form the open shells of the monomer’s triplet state. It can be seen, that already for the monomer, the spin-free orbital entropies are invariant to the choice of the $M_{s}$ component of the triplet state, while the “traditional” entropies change their values. In particular, the spin coupling of the active orbitals dramatically increases their entropy with respect to the spin-free one. Since we employed a determinant-based, not spin-adapted FCI code, for the dimer it did not always converge to an eigenfunction of $S^{2}$, but to an arbitrary linear combination of the degenerate eigenfunctions. By using a random initial guess, we achieved a fractional value of $\langle S^{2}\rangle$, yet the spin-free entropies were not affected. Moreover, for a spin-unpure state of the dimer the “traditional” entropies can differ for the two degenerate orbitals belonging to individual monomers, while the spin-free ones are the same.
@@
Orbital entropies
System $\langle S^{2}\rangle$, $M_{s}$ DOCC 1 DOCC 2 ACTIVE 1 ACTIVE 2 VIRT 1 VIRT 2
@@
Figure 2: Valence molecular orbitals of the CH2 diradical.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DOCC 1 DOCC 2 ACTIVE 1 ACTIVE 2 VIRT 1 VIRT 2
For the mutual information, there are too many data to show them all in a table, but it again turns out that the spin-free mutual information is invariant to the choice of spin state of the noninteracting dimer and is (up to a permutation of degenerate orbitals) a direct sum of the mutual information matrices of the monomers, while the “traditional” mutual information does not have these properties. The most interesting is the 4 x 4 submatrix corresponding to the two pairs of open shell orbitals of the monomers. The spin-free mutual information in this subspace vanishes, while the “traditional” one can acquire large values up to 0.31 for some spin couplings. It is thus evident that the spin-free correlation analysis has some advantages, being useful at least as a supplementary to the “traditional” one.
@@ 4.2 A complex with one iron atom
As a first realistic example of the presented spin-free correlation analysis we turn to the [Fe(SCH3)4]- complex, which contains an Fe(III) center in a S4 symmetry coordination environment. The ground state follows the Hund’s rule and turns to be a sextet, with each of the five MOs formed from 3d-orbitals of iron being singly occupied. The DMRG calculations of the sextet for the three values of $M_{s}$ = 1/2, 3/2, and 5/2 yield energy -3011.13544(6-8) a.u., with differences in the order of micro Hartrees. Reconstruction of the leading terms of the wave function from the MPS form confirmed that the $M_{s}=5/2$ component is single-determinantal, while the $M_{s}=3/2$ component is dominated by 5 equally contributing Slater determinants, where four of the MOs corresponding to the iron 3d-orbitals have $\alpha$-spin and one has $\beta$ spin. For $M_{s}=1/2$, (${5\choose 2}=10$ Slater determinants with three $\alpha$ and two $\beta$ spins contribute equally.
Figure 3: Correlation analysis for the ground state ($S=5/2$) of the [Fe(SCH3)4]- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well. The top row contains results for $M_{s}=5/2$ , middle row corresponds to $M_{s}=3/2$ and the bottom one has $M_{s}=1/2$. In the correlation graph, the orbitals in their ascending energy order are placed in the clockwise direction starting from the gray indicated “noon”. Configuration-averaged ROHF orbitals have been employed.
@@
Figure 3 shows the correlation analysis for this state and its three positive $M_{s}$ components, in the spin-dependent and spin-free forms. It can be seen that the three spin-free subfigures in the right column are identical, demonstrating the invariance of the spin-free entropy and mutual information with respect to the $M_{s}$ value. The top left subfigure corresponds to the spin-dependent analysis for the $M_{s}=5/2$ component, which is dominated by a single Slater determinant, and thus yield virtually identical results to the spin-free analysis. On the other hand, the middle and bottom left subfigures for $M_{s}=3/2$ and $M_{s}=1/2$, respectively, show strong orbital entropies for the MOs formed from 3d-orbitals of iron, and strong mutual information between each pair of them. For $M_{s}=1/2$ the values are somewhat bigger, in line with 10 Slater determinants contributing rather than five.
@@
The approach can be used for excited states as well. For example, the first excited state with energy -3011.0279(49-50) a.u. is a quartet, with its $M_{s}=3/2$ component dominated by the determinant $2\alpha\alpha\alpha 0$ with a non-negligible contribution of $0\alpha\alpha\alpha 2$, while the $M_{s}=1/2$ component has equal dominant contribution of determinants $2\beta\alpha\alpha 0$, $2\alpha\beta\alpha 0$, and $2\alpha\alpha\beta 0$ with a non-negligible contribution of analogous $0\ldots 2$ determinants. Figure 4 shows the corresponding correlation analysis for the two positive $M_{s}$ components of this state. It can again be seen that the spin-free entropies and mutual informations are invariant to $M_{s}$, and they agree with the high-spin spin-dependent case, showing strong entropies and mutual information for the first and last of the iron-3d dominated MOs. The correlation analysis for the $M_{s}=1/2$ component in the spin-dependent form shows in addition to this strong entropies and mutual information between the remaining three iron-3d dominated MOs, which are purely a consequence of the spin coupling.
@@
Figure 5 shows the corresponding correlation analysis for the second excited state, which is a doublet with energy -3011.016680 a.u.. Its wave function is dominated by several triples of equally contributing Slater determinants, where one of the 5 iron 3d MOs is doubly occupied, one empty, and the remaining three have two $\alpha$ and one $\beta$ electrons, in all three possible orderings. The static correlation due to spin couplings is thus inherently mixed with the strong non-dynamic correlation. This is reflected in the graphical analysis by the simplification of the mutual information graph when going from the spin-dependent to the spin-free analysis, and somewhat lower spin-free orbital entropies, particularly for one of the five orbitals.
@@
Table 2: Spin-free and spin-including total quantum informations $\tilde{S}_{\rm tot}$, $S_{\rm tot}$ and mutual informations $\tilde{I}_{\rm tot}$, $I_{\rm tot}$ for three states of the [Fe(SCH3)4]- complex and their $M_{s}$ components, computed using configuration averaged ROHF orbitals.
@@
5/2 5/2 0.085 0.087 0.044 0.082
5/2 3/2 0.085 2.588 0.044 3.354
5/2 1/2 0.085 3.451 0.044 4.556
3/2 3/2 0.879 0.891 0.898 1.082
3/2 1/2 0.879 2.792 0.898 2.960
1/2 1/2 3.709 5.623 2.208 4.372
An overview of the spin-free and spin-including total quantum informations and mutual informations for the three states of this complex are given in Tab. 2. As can be seen, the spin-free quantities are invariant with respect to $M_{s}$, while the spin-including ones strongly depend on the spin projection, with a growing trend towards smaller $M_{s}$. This can be explained by the increased number of Slater determinants forming low-$M_{s}$ components of the spin eigenfunctions, resulting in increased static correlation. However, even the spin-free total quantum information and mutual information have a growing trend with decreasing spin $S$, indicating more correlation present in the lower-spin states. This trend is somewhat hidden in the spin-including correlation measures due to their fluctuation with $M_{s}$, so in our view the spin-free quantities are a useful tool to uncover this dependence.
Table 3: Spin-free and spin-including total quantum informations $\tilde{S}_{\rm tot}$, $S_{\rm tot}$ and mutual informations $\tilde{I}_{\rm tot}$, $I_{\rm tot}$ for three states of the [Fe(SCH3)4]- complex and their $M_{s}$ components, computed using ROHF orbitals optimized for different spin states.
$S=5/2$ orbitals $S=3/2$ orbitals $S=1/2$ orbitals
$S$ $M_{s}$ $\tilde{S}_{\rm tot}$ $S_{\rm tot}$ $\tilde{I}_{\rm tot}$ $I_{\rm tot}$ $\tilde{S}_{\rm tot}$ $S_{\rm tot}$ $\tilde{I}_{\rm tot}$ $I_{\rm tot}$ $\tilde{S}_{\rm tot}$ $S_{\rm tot}$ $\tilde{I}_{\rm tot}$ $I_{\rm tot}$
5/2 5/2 0.029 0.031 0.025 0.031 0.055 0.057 0.032 0.052 0.119 0.121 0.059 0.113
5/2 3/2 0.029 2.532 0.025 3.303 0.055 2.557 0.032 3.325 0.119 2.622 0.059 3.385
5/2 1/2 0.029 3.395 0.025 4.506 0.055 3.420 0.032 4.527 0.119 3.484 0.059 4.588
3/2 3/2 0.892 0.911 0.859 1.066 0.854 0.866 0.881 1.054 0.913 0.925 0.919 1.118
3/2 1/2 0.892 2.811 0.859 2.938 0.854 2.767 0.881 2.933 0.913 2.826 0.919 2.993
1/2 1/2 3.662 5.584 3.281 5.694 3.617 5.532 2.487 5.163 3.740 5.653 2.938 5.390
[…]
We have thus seen that for states of all spin multiplicities, the spin-free correlation analysis was able to simplify the picture by filtering out the static correlation due to spin couplings. In particular, the spin-dependent mutual information graph looks almost identical for the $M_{s}=1/2$ component of the doublet and sextet states, while their spin-free graphs are dramatically different, reflecting the different origin of the correlation. It thus seems to be useful to accompany the “traditional” correlation analysis with its spin-free version.
Figure 4: Correlation analysis for the first excited state ($S=3/2$) of the [Fe(SCH3)4]- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well. The top row contains results for $M_{s}=3/2$, while the bottom one has $M_{s}=1/2$. Configuration averaged ROHF orbitals have been employed.
@@
Figure 5: Correlation analysis for the second excited state ($S=1/2$) of the [Fe(SCH3)4]- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well. Configuration averaged ROHF orbitals have been employed.
@@ 4.3 A complex with two iron atoms
The [Fe2S2(SCH3)4]2- complex is a model of the biologically important iron-sulfur clusters, which act as catalytic centers [34]. This species exhibits strong electronic correlation due to the coupling of the 3d-electrons of the two iron atoms, while the dynamic correlation is very important to get the correct energetic order of the states of different spin multiplicities (i.e. singlet ground state) [29]. In this work we did not aim to perform high-accuracy calculations of the energies of the states, rather we used the complex as a non-trivial example for the correlation analysis. Table 4 shows the energies of the $S=0-5$ states if this complex, computed in two different DMRG active spaces. As can be seen, DMRG with a very small active space like e.g. (14,14) yields inverted energy order of the states, however, DMRG(22,21) already yields a qualitatively correct ordering, although the excitation energies are substantially underestimated with respect to DMRG-AC0 [29] due to the lack of dynamic correlation.
@@
Figure 6: Correlation analysis for the ground ($S=0$) state of the [Fe2S2(SCH3)4]2- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well.
@@
Fig. 6 shows the correlation analysis for the ground ($S=0$) state. It clearly shows the presence of the full pairwise “anti-ferromagnetic” coupling between the pairs formed from the 3d-dominated MOs of the two iron atoms. Smaller contributions in the wave function come from determinants with some open shells, which have to spin-couple to singlets, and this is reflected by the weak mutual information lines, which disappear in the spin-free subfigure on the right, and the slightly lower spin-free orbital entropies.
Figure 7: Correlation analysis for the first excited ($S=1$) state of the [Fe2S2(SCH3)4]2- complex, spin-including (left) and spin-free (right). The bar graphs shows orbital entropies, while the color-coded size of the mutual information matrix elements is displayed below. The weighted graphs combining the orbital entropies and mutual information are plotted as well. The $M_{s}=1$ component is in the top row, with $M_{s}=0$ component below.
@@
Fig. 7 shows the results for the first excited ($S=1$) state in its two $M_{s}=1,0$ components. The overall situation is similar to the ground state, except that here one pair of orbitals remains open shell “ferromagnetically” triplet-coupled. The spin-including mutual informations differ for the two $M_{s}$ components, while the spin-free ones are identical and a similar “weakening” of the mutual information and orbital entropies as in the ground state happens. The situation is analogous for the other excited states, with the most dramatic spin-free “simplification” occurring for the $M_{s}<S$ components of the “totally ferromagnetic” $S=5$ excited state (analogous to the $S=5/2$ state of the single-iron complex, cf. Fig. 3). Rather than showing figures for all the cases, we summarize the results in Tab. 5. The somewhat less than perfect $M_{s}$-invariance of the spin-free quantities for higher DMRG roots corresponds to the truncation of the bond dimension, which was limited to 2048 for all states and is more limiting to the low-$M_{s}$ states being higher roots. In some cases we did not achieve convergence at all for the higher roots. Nevertheless, the total spin-free entropy $\tilde{S}_{\rm tot}$ shows the same increasing trend towards low spin states as in the complex with single iron atom. On the other hand, the total spin-free mutual information $\tilde{I}_{\rm tot}$ is very similar (around 2.2) for states with $S=0-4$ and decreases to 0.13 only for the highest spin state $S=5$. This is in line with the observation from Figs. 6,7, which show strong pairwise correlation between orbitals from the two iron atoms, both for the ground state $(S=0)$ and first excited state $(S=1)$. This correlation “survives” in the spin-free analysis and is thus not caused by spin coupling, but reflects the inherent strong correlation in the system.
Table 5: Spin-free and spin-including total quantum informations $\tilde{S}_{\rm tot}$, $S_{\rm tot}$ and mutual informations $\tilde{I}_{\rm tot}$, $I_{\rm tot}$ for three states of the [Fe2S2(SCH3)4]2- complex and their $M_{s}$ components.